OpenStudy (konradzuse):

Integration by parts q. e^2x sin(3x) dx

OpenStudy (konradzuse):

I know how to do it, but in this situation when I get to e^2x sin(3x) dx = something integral e^2x sin(3x) dx do I divide or do I - out one from the other..?

OpenStudy (konradzuse):

There's some special rule about this, I forgot how to do it :).

OpenStudy (amistre64):

running thru twice tends to get it you to a point where you can algebra it into submission

OpenStudy (lgbasallote):

\[\huge \int e^{2x} \sin (3x) dx?\]

OpenStudy (konradzuse):

yeah iggy.

OpenStudy (amistre64):

since e^2x is not integratable; use it as the u parts

OpenStudy (konradzuse):

that's only because we don't have a du if we were to say u = 2x right?

OpenStudy (konradzuse):

I guess I'll just do the problem here then.

OpenStudy (amistre64):

correct

OpenStudy (lgbasallote):

but according to LIATE you should use sin (3x) right @amistre64 ?

OpenStudy (amistre64):

according to common sense .... LIATE is jsut suggestions that are common, but not always applicable

OpenStudy (konradzuse):

u = 2x 1/2 du = dx dv = sin(3x) v = 3cos(3x)?

OpenStudy (lgbasallote):

u = 2x?

OpenStudy (amistre64):

u = e^2x du = 2e^2x dx dv = sin(3x) v = -cos(3x)/3

OpenStudy (konradzuse):

actually I guess I gotta say u = e^2x.

OpenStudy (lgbasallote):

shouldnt it be e^2x?

OpenStudy (konradzuse):

Yeah sorry I messed up stupidly.

OpenStudy (konradzuse):

couple times.

OpenStudy (amistre64):

youll have to run it thru a second time, and you should get the same integral as you started with, or some multiplied version of it

OpenStudy (konradzuse):

yuh... Do I divide that?

OpenStudy (konradzuse):

I'm going to do this on paper, I make more mistakes when I type everything fast on here.

OpenStudy (amistre64):

\[\int udv = u_1v_1-\int v_1du_1\] \[\int udv = u_1v_1-(v_2u_2 -\int u_2dv_2)\] \[\int udv = u_1v_1-v_2u_2 +\int u_2dv_2\] \[A = u_1v_1-v_2u_2 +A\]

OpenStudy (konradzuse):

so you could subtract it then?

OpenStudy (amistre64):

yes, you would subtract it; combine it with the other side; and divide off the "coefficient"

OpenStudy (amistre64):

itll look better after you work it to that step and go ... ohhh!!

OpenStudy (konradzuse):

if it's exactly the same wouldn't it then be = 0?

OpenStudy (amistre64):

IF, then yes; but its not going to be exactly the same, it was just easier to notate it inthe latex like that for simplicity

OpenStudy (konradzuse):

alright so fa rI have

OpenStudy (amistre64):

\[A = u_1v_1-v_2u_2 +nA\] \[A-nA = u_1v_1-v_2u_2\] \[(1-n)A = u_1v_1-v_2u_2\] \[A = \frac{u_1v_1-v_2u_2}{1-n}\]

OpenStudy (konradzuse):

\[\frac{e^{2x} \cos(3x)}{3} + 1/3 \int\limits \cos(3x) 2 e^{2x}dx\]

OpenStudy (konradzuse):

I could throw the 2 out in front right?

OpenStudy (amistre64):

\[\int e^{2x} sin(3x) dx= -\frac{e^{2x} \cos(3x)}{3} +\frac 23 \int\limits e^{2x} cos(3x) dx\] the 2s a constant that can be pulled out yes

OpenStudy (konradzuse):

u = e^2x du = 2e^2x dx

OpenStudy (konradzuse):

dv = cos(3x)dx v = 1/3 sin(3x)

OpenStudy (amistre64):

i cant see the preview so i hope this is right :) \[\int e^{2x} sin(3x) dx= -\frac{e^{2x} \cos(3x)}{3} +\frac 23 \left(\frac{e^{2x}sin(3x)}{3}-\frac 23 \int\limits e^{2x} sin(3x) dx\right)\]

OpenStudy (amistre64):

yes

OpenStudy (konradzuse):

so it's multiplied when you do it a second time? I tohught you just add or subtract?

OpenStudy (amistre64):

no, your simply doing the int by parts again but its all multiplied by the outer constant, -2/3 in this case

OpenStudy (konradzuse):

OH IC....

OpenStudy (konradzuse):

yeah the original 2/3 gotcha... I thought it was multiplying by the first integration by parts we did I was like wuh... :)

OpenStudy (konradzuse):

so now we are stuck with 2/3 sin(3x)e^2xdx

OpenStudy (amistre64):

\[\int A dx= -\frac{e^{2x} \cos(3x)}{3} +\frac 23 \left(\frac{e^{2x}sin(3x)}{3}-\frac 23 \int A dx\right)\] stuck with? not quite \[\int A dx= -\frac{e^{2x} \cos(3x)}{3} +\frac 23 \left(-\frac 29 {e^{2x}sin(3x)}-\frac 49 \int A dx\right)\]

OpenStudy (konradzuse):

oh woops I did 4/6 for some reason LOL.

OpenStudy (amistre64):

\[\int A dx= -\frac{e^{2x} \cos(3x)}{3} +-\frac 29 {e^{2x}sin(3x)}-\frac 49 \int A dx\]

OpenStudy (konradzuse):

which would go back to 2/3... Man I need to get something to eat soon no brain food...

OpenStudy (amistre64):

i got an extra "-" lol :)

OpenStudy (amistre64):

as you can see, hopefully; we can algebra the rest

OpenStudy (amistre64):

\[\int A dx= -\frac 13 {e^{2x} \cos(3x)} +\frac 29 {e^{2x}sin(3x)}-\frac 49 \int A dx\] \[\int A dx+\frac 49 \int A dx= -\frac 13 {e^{2x} \cos(3x)} +\frac 29 {e^{2x}sin(3x)}\] \[\int A dx(1+\frac 49)= -\frac 13 {e^{2x} \cos(3x)} +\frac 29 {e^{2x}sin(3x)}\] \[\int A dx= \frac{-\frac 13 {e^{2x} \cos(3x)} +\frac 29 {e^{2x}sin(3x)}}{1+\frac 49}\]

OpenStudy (konradzuse):

sorry I had to help my parents with the groceries and such...

OpenStudy (konradzuse):

so could we combine the 9/9 + 4/9 to get 13/9 then so a flip?

OpenStudy (konradzuse):

@amistre64

OpenStudy (turingtest):

what do you mean a "flip" ?

OpenStudy (turingtest):

oh yeah, I see what you mean yeah, flip it

OpenStudy (konradzuse):

Alrighty, gotta feed my dog's I'll be back, sorry.

OpenStudy (konradzuse):

Alrighty, gotta feed my dog's I'll be back, sorry.

OpenStudy (konradzuse):

@amistre64 I was curious how it turned into +2/9sin(3x)?

OpenStudy (konradzuse):

\[\int\limits A dx= -\frac{e^{2x} \cos(3x)}{3} +\frac 23 \left(-\frac 29 {e^{2x}\sin(3x)}-\frac 49 \int\limits A dx\right)\]

OpenStudy (konradzuse):

Oh I think this was accidentally added, the one before wasn't -, and the one after isn't... That is where I was confused okay....

OpenStudy (konradzuse):

I got 3/13 e^2xcos(3x) + 3/26 e^2x sin(3x) = Adx

OpenStudy (konradzuse):

@TuringTest